OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−4, −1) and (4, 5)?

OpenStudy (anonymous):

@jim_thompson5910 @Directrix

OpenStudy (anonymous):

@cwrw238 @radar @zepdrix

OpenStudy (anonymous):

change it to slope intercept equation y=mx+b b=y-int m=slope

OpenStudy (anonymous):

to calculate for slope: \[m=\frac{ \Delta y }{ \Delta x }=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

using the given coordinates calculate for b

OpenStudy (anonymous):

So I substitute the variables for the coordinates?

OpenStudy (anonymous):

kinda...?

OpenStudy (anonymous):

\[\frac{ -1 - 5 }{ -4 - 4 }\]

OpenStudy (anonymous):

Like that?

OpenStudy (anonymous):